aircraft travels 864 miles with the wind in 4 hours and the same distance against the wind in 6 hours. What is the speed of the wind?
Answer by mananth(16949)

aircraft travels 864 miles with the wind in 4 hours and the same distance against the wind in 6 hours. What is the speed of the wind?
...
plane speed in still air =x
wind speed = y
...
with wind speed = x+y
against wind speed = x-y
...
t=d/r
864/(x+y)=4
864=4x+4y
/4
216=x+y....................1
..
864/(x-y)=6
6x-6y=864
x-y=144.......................2
add both equations
..
2x=360
x= 180 mph speed of plane
y = 36 speed of wind
...
m.ananth@hotmail.ca